OpenStudy (anonymous):

Circle P has a radius of 8 units with center P at (2, -1). Which equation defines circle P?

OpenStudy (anonymous):

(x-2)^2 + (y+1)^2 = 64

OpenStudy (anonymous):

can you explain how you got that please?

OpenStudy (shaik0124):

(x-2)^2+(y+1)^2=8^2

OpenStudy (shaik0124):

the eeauation of circle with 1 point(x,y) and radius r is given by (x-x1)^2+(y-y1)^2=r^2

OpenStudy (shaik0124):

substitute x1 and y1 in the circle equation u will the answer

OpenStudy (anonymous):

y an x are the center, the radius is the square root of the of 64.

OpenStudy (anonymous):

Thank you!!!